# Full private delegated quantum computing tailored from user to industry

### v1

@@ Abstract
Quantum computing, Delegated quantum computing, Quantum cryptography, Circuit-based blind quantum computation
preprint: APS/123-QED
I Introduction
@@
The structure of this paper is as follows. First in Sec. II we will begin by explaining briefly how the main techniques known in the field work, and then we will propose combinations and modifications of these techniques to adapt them to the needs and requirements of the client, which may be a company in Sec. III or a user with fewer resources in Sec. IV. Finally, in Sec. VI we will give 3 examples of how such protocols would be realized in a particular algorithm.
II Delegated quantum computing techniques
@@ II.3 FBQC
III Protocols adapted to industry needs
In this section we are going to analyze the different cases that can occur in the context of a company as a client, with larger quantum resources and problem sizes. First, in Ssec. III.1 we will study the unconstrained case, only with a memory size problem, which will serve as an initial scheme to understand the basis of the protocols we will introduce. After that, in Ssec. III.2 we will address the case of having a constraint on the possible client operations. Specifically, the client will not be able to perform multi-qubit operations, in other words, instead of having one quantum computer of many qubits, the client will have many quantum computers of one qubit.
III.1 $M$ qubits total computer
@@
However, we want the data embedded in the states we send back to be completely hidden from the external server, so we perform one-time pad encryption like explained in Ssec. II.2. This way, when we are going to send a qubit to the external server, we will only have to apply an $X^{a}Z^{b}$ operation to the qubit in question with the key $(a,b)$ and then undo it when we receive it back, using the methods of [29, 15, 16, 11]. For the generation of the random keys for encryption, one option is dedicate a qubit to generate the $a$ and $b$. Other one is to generate all the necessary keys prior to the execution of the algorithm and thus be able to perform the protocol with all the qubits. However, this is not possible in all algorithms, such as in an HHL algorithm [30], since we do not know in advance how many repetitions of the circuit we will have to perform, and therefore, the length of the final circuit. The third option is to use an external classical random number generator, which may compromise the randomness of the keys. The fourth option, and best one, will consist of the Protocol 1 every time we have free slots in our computer after sending a round of qubits to the external server.
@@
III.2 $M$ one-qubit total computers
@@
IV Protocols adapted to user requirements
@@
IV.1 $2$ qubits total computer
The case of a 2-qubit device that can perform all possible operations with them is a particular case of the problem with $M$ qubits of Sec. III.1, so the only modification to be made is to impose that $M=2$.
@@ IV.1 $2$ qubits total computer
IV.2 $2$ one-qubit total computers
@@
IV.3 $1$ qubit total computer
@@
However, a very important problem arises: we cannot use the $SWAP$s method to disguise the structure of the client’s operations. Therefore, in cases where we need to cover up the structure of the operations, we will have to rely only on being able to introduce unknown gates for the server or use an MBQC method, with the cost in server resources that this implies. We can also try to combine the Ssec. IV.5 technique with the client qubit to deal with the $SWAP$s problem by making it an external node.
IV.4 $M$ one-qubit partial computers
@@ V Verification algorithm
VI Examples with quantum algorithms

### v2

@@ Abstract
1 Introduction
@@
The structure of this paper is as follows. First in Sec. 2 we will begin by explaining briefly how the main techniques known in the field work, and then we will propose combinations and modifications of these techniques to adapt them to the needs and requirements of the client, which may be a company in Sec. 3 or a user with fewer resources in Sec. 4. Finally, in Sec. 6 we will give 3 examples of how such protocols would be realized in a particular algorithm.
2 Delegated quantum computing techniques
@@ 2.3 FBQC
3 Protocols adapted to industry needs
In this section we are going to analyze the different cases that can occur in the context of a company as a client, with larger quantum resources and problem sizes. First, in Ssec. 3.1 we will study the unconstrained case, only with a memory size problem, which will serve as an initial scheme to understand the basis of the protocols we will introduce. After that, in Ssec. 3.2 we will address the case of having a constraint on the possible client operations. Specifically, the client will not be able to perform multi-qubit operations, in other words, instead of having one quantum computer of many qubits, the client will have many quantum computers of one qubit.
3.1 $M$ qubits total computer
@@
However, we want the data embedded in the states we send back to be completely hidden from the external server, so we perform one-time pad encryption like explained in Ssec. 2.2. This way, when we are going to send a qubit to the external server, we will only have to apply an $X^{a}Z^{b}$ operation to the qubit in question with the key $(a,b)$ and then undo it when we receive it back, using the methods of [29, 15, 16, 11]. For the generation of the random keys for encryption, one option is dedicate a qubit to generate the $a$ and $b$. Other one is to generate all the necessary keys prior to the execution of the algorithm and thus be able to perform the protocol with all the qubits. However, this is not possible in all algorithms, such as in an HHL algorithm [30], since we do not know in advance how many repetitions of the circuit we will have to perform, and therefore, the length of the final circuit. The third option is to use an external classical random number generator, which may compromise the randomness of the keys. The fourth option, and best one, will consist of the Protocol 1 every time we have free slots in our computer after sending a round of qubits to the external server.
@@
3.2 $M$ one-qubit total computers
@@
4 Protocols adapted to user requirements
@@
4.1 $2$ qubits total computer
The case of a 2-qubit device that can perform all possible operations with them is a particular case of the problem with $M$ qubits of Sec. 3.1, so the only modification to be made is to impose that $M=2$.
@@ 4.1 $2$ qubits total computer
4.2 $2$ one-qubit total computers
@@
4.3 $1$ qubit total computer
@@
However, a very important problem arises: we cannot use the $SWAP$s method to disguise the structure of the client’s operations. Therefore, in cases where we need to cover up the structure of the operations, we will have to rely only on being able to introduce unknown gates for the server or use an MBQC method, with the cost in server resources that this implies. We can also try to combine the Ssec. 4.5 technique with the client qubit to deal with the $SWAP$s problem by making it an external node.
4.4 $M$ one-qubit partial computers
@@ 5 Verification algorithm
6 Examples with quantum algorithms